OpenStudy (anonymous):

Please Help! Given the system of equations, match the following items. 2x - y = 0 x + y= -3

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & -1 \\ -3 & 1\end{matrix}\right] \left[\begin{matrix}2 & 0 \\ 1 & -3\end{matrix}\right] \left[\begin{matrix}2 & -1 \\ 1 & 1\end{matrix}\right]\] x-determinant y-determinant System determinant

OpenStudy (anonymous):

@texaschic101 @satellite73

OpenStudy (anonymous):

\[16 \sqrt{9^5} - 8 \sqrt{a^3} + 2 \sqrt{a} \over 2\sqrt{a}\]